From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Mar 19 2003 - 23:49:51 EST
------------------------------------------------------------ Date: Wed Mar 19 23:49:51 2003, Shift: 16:00-24:00 Supervisor: Michael, with: Log book#14 pages 68 - 69 Shift Summary: ------------------------------------------------------------ 1) After access beam came back around 9pm and has been on since then. TOFW/ZDC has been about 10. ZDC scaler was 7000 now about 3600. 2) Before beam we took a set of pedestal runs. Since 9pm we have completed the FS 1/4A running (203K in run 8514). Currently run 8515 has 237k at 1/8A. 3) T5 Y2 seemed to lose one plane. Three channels in TOF2 have high voltage set to zero. On DAQ the mtpc1-comm panel has flashed Red 2 or 3 times this shift. At 9pm I noticed that all the magnets had fallen into "standby -fault". This has not happened since. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Wed Mar 19 2003 - 23:50:55 EST